OpenStudy (anonymous):

Find the series solution (centered at zero) to the differential equation y'+12xy=0. Find the general term of the series and identify it as the sum, product, quotient, or composition of the library of algebraic and/or transcendental functions.

OpenStudy (anonymous):

To get an idea of the solution, notice that your DE is a separable on \[y'+12xy=0\\ \frac {y'}y=-12 x\\ \]

OpenStudy (anonymous):

\[ \ln(y/c) = -6 x^2\\ y= c e^{-6 x^2} \]

OpenStudy (anonymous):

Your series solutions, also will give you that. Try to do it.

OpenStudy (anonymous):

Your series should look like \[c-6 c x^2+18 c x^4-36 c x^6+54 c x^8-\\ \frac{324 c x^{10}}{5}+O\left(x^{11}\right ) \]

OpenStudy (anonymous):

Would i start with